OpenStudy (anonymous):

One of the solutions to x2 − 2x − 15 = 0 is x = −3. What is the other solution? x = −5 x = −1 x = 1 x = 5

OpenStudy (anonymous):

The only solution of the equation x2 + bx + 16 = 0 is x = 4. What is the value of b? b = –16 b = –8 b = 8 b = 16

OpenStudy (anonymous):

An equation has solutions of m = –5 and m = 9. Which could be the equation? (m + 5)(m – 9) = 0 (m – 5)(m + 9) = 0 m2 – 5m + 9 = 0 m2 + 5m – 9 = 0

OpenStudy (anonymous):

Brianna is graphing the function f(x) = x2 + 6x + 5. What x-intercepts should Brianna use to graph f(x)? –5 and –1 –5 and 1 –1 and 5 1 and 5

OpenStudy (anonymous):

The function h(x) is quadratic and h(3) = h(–10) = 0. Which could represent h(x)? h(x) = x2 – 13x – 30 h(x) = x2 – 7x – 30 h(x) = 2x2 + 26x – 60 h(x) = 2x2 + 14x – 60

OpenStudy (anonymous):

The function g(x) = x2 – 10x + 24 is graphed on a coordinate plane. Where will the function cross the x-axis? (–6, 0) and (–4, 0) (–2, 0) and (–12, 0) (4, 0) and (6, 0) (12, 0) and (2, 0)

OpenStudy (susanna):

x=5

OpenStudy (susanna):

The only solution of the equation x2 + bx + 16 = 0 is x = 4. What is the value of b? Since the only solution is x=4, you can set up the factored version. (x-4)(x-4) Use FOIL. x^2-8x+16 So, b=-8

OpenStudy (susanna):

An equation has solutions of m = –5 and m = 9. Which could be the equation? You are given the solutions, now set the up in factored for. (m+5)(m-9) Because m+5=0 Subtract 5 from both sides ---> m=-5 and m-9=0 Add 9 to both sides ---> m=9

OpenStudy (susanna):

Brianna is graphing the function f(x) = x2 + 6x + 5. What x-intercepts should Brianna use to graph f(x)? -5 and -1 Graph it and see for yourself.

OpenStudy (susanna):

The function g(x) = x2 – 10x + 24 is graphed on a coordinate plane. Where will the function cross the x-axis? (4, 0) and (6, 0) Graph it, and you'll see for yourself.

OpenStudy (susanna):

The function h(x) is quadratic and h(3) = h(–10) = 0. Which could represent h(x)? h(x) = 2x^2 + 14x – 60 Because h(3)=2(3)^2+14(3)-60 =18+42-60 =0

OpenStudy (susanna):

PLEASE post ONE question at a time!!

OpenStudy (anonymous):

thank youu